OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

tim is 13 and coach johnson is 23 now can you figure it out?

hero (hero):

She wants to understand the steps

OpenStudy (anonymous):

no

OpenStudy (bahrom7893):

C = 10+T.. right?

hero (hero):

Of course, I posted the steps in your previous question, but you deleted it again, even after I told you not to.

OpenStudy (anonymous):

I didn't

OpenStudy (anonymous):

Please help

hero (hero):

j = t + 10 j - 3 = 2(t-3) j = t + 10 j = 2t - 3 j = j t + 10 = 2t - 3 2t - t = 10 + 3 t = 13 j = 23 Three years ago, t = 10, j = 20

OpenStudy (anonymous):

So how old is tim now?

hero (hero):

t = 13

hero (hero):

t = Tim

hero (hero):

Tim is 13

hero (hero):

I just said it three different ways

OpenStudy (mertsj):

Way to go, Hero!!! you're a legend!!

hero (hero):

You're just now finding that out? But thanks anyways